Rot KE = \frac{1}{2}Iω^{2}

Rot KE_{f} = Rot KE_{i} + Work_{net}
Rot KE_{f} = Rot KE_{i} + ∫ τ⃗_{net} · d⃗θ

W = ∫ τ⃗ · d⃗θ

power P = \frac{dW}{dt} = \frac{d}{dt} ∫ τ⃗ · d⃗θ

= τ⃗ · \frac{d⃗θ}{dt}

= τ⃗ · ω⃗ Work = (Power) (time)